OpenStudy (anonymous):

A system of equations is shown below: 5x + 2y = 3 (equation 1) 2x – 3y = 1 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof?

OpenStudy (anonymous):

A) Show that the solution to the system of equations 7x – y = 4 and 2x - 3y = 1 is the same as the solution to the given system of equations B) Show that the solution to the system of equations 2x + 5y = 3 and 3x - 2y = 1 is the same as the solution to the given system of equations C)Show that the solution to the system of equations 9x + 4y = 5 and 7x - y = 4 is the same as the solution to the given system of equations D) Show that the solution to the system of equations -4x + 9y = 5 and 2x - 3y = 1 is the same as the solution to the given system of equations

OpenStudy (anonymous):

help please :(

OpenStudy (mathmath333):

find the point of intersection of the lines \(\large 5x + 2y = 3~~ and~~ 2x – 3y = 1\)

OpenStudy (mathmath333):

point of intersection of the lines \(\large 5x + 2y = 3~~ and~~ 2x - 3y = 1\) will match one of the point of intersection of the two lines given in options ,try first option \(\large 7x – y = 4 ~~and ~~2x-3y = 1\)